OpenStudy (anonymous):

A material decays at a rate of 1.1% per year. If you start with 250 grams of the material, how much will be left in 15 years? Round to the nearest whole gram. Which equation correctly models the situation?

OpenStudy (anonymous):

A. 250= A(.011)^15 B. 250= A(.989)^15 C. A= 250(.011)^15 D. A= 250(.989)^15

OpenStudy (anonymous):

@Hero @KendrickLamar2014 @ganeshie8 Please help me!!!(:

OpenStudy (kendricklamar2014):

211.7797

OpenStudy (anonymous):

Oh well solving it was the next question. So thank you!! Which equation is it? A B C or D?

OpenStudy (kendricklamar2014):

C